OpenStudy (anonymous):

Use Euler’s formula to rewrite this radical expression in exponential form. 2√2 - 2i√2

OpenStudy (anonymous):

I have Euler's Formula as:\[a+bi=r(\cos \Theta+i \sin \Theta)\] But how do i plug it in...??

OpenStudy (kc_kennylau):

This is not Euler's Formula?

OpenStudy (kc_kennylau):

I don't know much about this; but do you have formulas for the values of r and θ?

OpenStudy (anonymous):

My textbook says that's the formula for a complex number, a+bi. I also have \[e^(i \alpha) = \cos \alpha+i \sin \alpha \] but that makes even less sense. This is literally all the information i am given. that's why i'm confused.

OpenStudy (kc_kennylau):

Oh I understand now; you have to express \(2\sqrt2-2i\sqrt2\) in the form of \(r(\cos\theta+i\sin\theta)\) first.

OpenStudy (anonymous):

.... how do i do that? :(

OpenStudy (kc_kennylau):

To express \(a+bi\) in the form of \(r(\cos\theta+i\sin\theta)\): \[r=\sqrt{a^2+b^2}\]\[\tan\theta=\frac ba\]

OpenStudy (kc_kennylau):

Or you can solve equations: \[a=r\cos\theta\]\[bi=ri\sin\theta\]

OpenStudy (kc_kennylau):

And you'll come up with my formulas

zepdrix (zepdrix):

Go with Kenny's method, it's more reliable, but here's another idea! This is what I like to do at least :) \[\Large\rm \color{royalblue}{2\sqrt2} - \color{orangered}{2\sqrt2}\mathcal i\] Try to turn each term into the special values that correspond to the unit circle. Factoring a 4 out of each term,\[\Large\rm 4\left(\color{royalblue}{\frac{\sqrt2}{2}}-\color{orangered}{\frac{\sqrt2}{2}}\mathcal i\right)\] Then relating it to Euler,\[\Large\rm \color{royalblue}{\cos \theta=\frac{\sqrt2}{2}},\qquad\qquad\qquad \color{orangered}{\sin \theta=-\frac{\sqrt2}{2}}\]Sine is negative, cosine is positive, uhhh we must be in the 4th quadrant I guess. And we sqrt2's, so we must be dealing with one of the pi/4's right?

OpenStudy (anonymous):

Okay so i need to get this in exponential form... So, after i get it to \[4(\cos 7\Pi/4 + i \sin 7\Pi/4)\] how do i get it to exponential form???

OpenStudy (kc_kennylau):

\[\Large\cos\theta+i\sin\theta=e^{i\theta}\]

OpenStudy (anonymous):

... i dont understand x(

OpenStudy (kc_kennylau):

\(\theta=\dfrac{7\pi}4\)

OpenStudy (kc_kennylau):

\[\Large\cos\frac{7\pi}4+i\sin\frac{7\pi}4=e^{i\frac{7\pi}4}\]

OpenStudy (anonymous):

OOOHHH, I SEE! Hahaha wow, okay, yeah. Thank you soooo so much! :))))

OpenStudy (kc_kennylau):

no problem :)